# SurveyG: A Multi-Agent LLM Framework with Hierarchical Citation Graph for Automated Survey Generation

### v2

@@ B.1. Survey Topics
B.2. Evaluation Metrics
B.2.1. Metrics about Content Quality
We evaluate both the quality of the generated outlines and the full survey papers. A well-structured and logically coherent outline is essential for maintaining clarity and organization, and we adopt the same evaluation settings as in (Yan et al., 2025). The full paper evaluation serves as a comprehensive qualitative benchmark to assess the academic rigor and practical utility of the generated surveys. Following the prompt design and evaluation protocols from previous studies (Wang et al., [n. d.]; Liang et al., 2025; Yan et al., 2025; Su et al., 2025), we assess survey quality across five key metrics: Coverage, which measures how thoroughly the survey captures major concepts, foundational works, and emerging trends; Structure, which examines logical organization, coherence, and taxonomy quality; Relevance, which assesses the alignment of content with the target research topic; Synthesis, which evaluates the integration of information from multiple sources into a cohesive and non-redundant narrative; and Critical Analysis, which reflects the survey’s ability to identify methodological gaps, highlight trends, and articulate open research challenges. Each metric is scored on a 0–100 scale by both LLM-based judges and human experts, with higher scores indicating stronger performance. The complete evaluation prompts and scoring criteria are detailed in Appendix B.
B.2.2. Metric about Citation Quality
Following the methodology in prior studies (Wang et al., [n. d.]; Gao et al., 2023), we evaluate the citation quality of the generated surveys by measuring both the accuracy and the contextual relevance of cited references. Specifically, we extract a set of factual claims from each generated survey and verify whether these claims are appropriately supported by their corresponding references. To automate this process, we employ a Natural Language Inference (NLI) model that determines whether the content of each cited paper logically supports the associated claim. Based on this evaluation, we calculate two key metrics: Citation Recall, which reflects the proportion of claims that are correctly supported by valid references, and Citation Precision, which measures the proportion of cited references that truly substantiate the claims they are linked to. Together, these metrics provide a robust measure of how accurately and meaningfully the generated surveys integrate citations within their arguments.
Appendix C Prompt Templates
@@
C.1. Prompt to generate structured outline
We provide a short version of the prompt template (Figure 7) that instructs the model to construct a coherent, hierarchical outline that captures the logical flow of a literature review topic before detailed writing begins.prior to
@@ C.1. Prompt to generate structured outline
C.2. Prompt to evaluate structured outline
@@
C.3. Prompt to generate subsections
@@
Figure 9. Generate Subsection Prompt Short Version
C.4. Improve Section Quality
As shown in Figure 10, this prompt systematically assesses literature review sections across multiple dimensions such as content coverage, synthesis, and critical analysis while offering actionable feedback and retrieval suggestions for refinement.
@@ C.4. Improve Section Quality
Appendix D Case studies
@@
Figure 11. Case studies about the result of generated subsection.

### v3

@@ Appendix A Details about Traversal on Graph
Appendix B More ablation study results
LLM Model Backbone Coverage Structure Relevance Synthesis Critical Analysis
Claude SurveyForge GPT-4o-mini 81.8(± 2.9)81.8\,($\pm\,$2.9) 78.4(± 3.5)78.4\,($\pm\,$3.5) 89.1(± 2.1)89.1\,($\pm\,$2.1) 75.4(± 3.8)75.4\,($\pm\,$3.8) 70.3(± 4.2)70.3\,($\pm\,$4.2)
SurveyForge Gemini-2.5-Flash 82.5(± 3.1)82.5\,($\pm\,$3.1) 79.2(± 3.7)79.2\,($\pm\,$3.7) 89.8(± 2.3)89.8\,($\pm\,$2.3) 76.8(± 3.9)76.8\,($\pm\,$3.9) 72.1(± 4.3)72.1\,($\pm\,$4.3)
SurveyG GPT-4o-mini ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{88.1}}\,(\pm\,2.4)$ 87.9(± 2.7)87.9\,($\pm\,$2.7) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{93.6}}\,(\pm\,1.9)$ 80.2(± 3.2)80.2\,($\pm\,$3.2) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{77.3}}\,(\pm\,3.5)$
SurveyG Gemini-2.5-Flash 87.5,(± 2.6)87.5,($\pm\,$2.6) $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}88.6}\,(\pm\,2.5)}$ 93,2(± 2.0)93,2\,($\pm\,$2.0) $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}81.4}\,(\pm\,3.1)}$ 76.9(± 3.6)76.9\,($\pm\,$3.6)
GPT SurveyForge GPT-4o-mini 94.2(± 1.8)94.2\,($\pm\,$1.8) 87.3(± 2.5)87.3\,($\pm\,$2.5) 94.8(± 1.6)94.8\,($\pm\,$1.6) 88.6(± 2.7)88.6\,($\pm\,$2.7) 88.5(± 2.9)88.5\,($\pm\,$2.9)
SurveyForge Gemini-2.5-Flash 94.8(± 1.9)94.8\,($\pm\,$1.9) 87.9(± 2.6)87.9\,($\pm\,$2.6) 95.2(± 1.5)95.2\,($\pm\,$1.5) 89.4(± 2.8)89.4\,($\pm\,$2.8) 89.3(± 3.0)89.3\,($\pm\,$3.0)
SurveyG GPT-4o-mini 95.7(± 1.5)95.7\,($\pm\,$1.5) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{88.5}}\,(\pm\,2.2)$ 95.1(± 1.4)95.1\,($\pm\,$1.4) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{92.2}}\,(\pm\,2.3)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{91.2}}\,(\pm\,2.6)$
SurveyG Gemini-2.5-Flash $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}96.3}\,(\pm\,1.4)}$ 88.2,(± 2.4)88.2,($\pm\,$2.4) $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}95.7}\,(\pm\,1.3)}$ 91.8(± 2.5)91.8\,($\pm\,$2.5) 90.5(± 2.5)90.5\,($\pm\,$2.5)
Deepseek SurveyForge GPT-4o-mini 89.4(± 2.3)89.4\,($\pm\,$2.3) 86.5(± 2.9)86.5\,($\pm\,$2.9) 92.5(± 2.0)92.5\,($\pm\,$2.0) 84.1(± 3.4)84.1\,($\pm\,$3.4) 80.7(± 3.7)80.7\,($\pm\,$3.7)
SurveyForge Gemini-2.5-Flash 88.9(± 2.5)88.9\,($\pm\,$2.5) 85.8(± 3.1)85.8\,($\pm\,$3.1) 93.1(± 2.1)93.1\,($\pm\,$2.1) 85.3(± 3.5)85.3\,($\pm\,$3.5) 81.5(± 3.8)81.5\,($\pm\,$3.8)
SurveyG GPT-4o-mini 88.7(± 2.5)88.7\,($\pm\,$2.5) 85.7(± 3.0)85.7\,($\pm\,$3.0) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{94.2}}\,(\pm\,1.8)$ 86.3(± 3.1)86.3\,($\pm\,$3.1) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{82.7}}\,(\pm\,3.5)$
SurveyG Gemini-2.5-Flash $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}89.1}\,(\pm\,2.4)}$ $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}86.1}\,(\pm\,2.8)}$ 93.8(± 1.9)93.8\,($\pm\,$1.9) $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}86.8}\,(\pm\,3.0)}$ 82.3(± 3.7)82.3\,($\pm\,$3.7)
Gemini SurveyForge GPT-4o-mini 94.9(± 1.7)94.9\,($\pm\,$1.7) 93.1(± 2.1)93.1\,($\pm\,$2.1) 98.7(± 1.1)98.7\,($\pm\,$1.1) 93.5(± 2.4)93.5\,($\pm\,$2.4) 94.2(± 2.5)94.2\,($\pm\,$2.5)
SurveyForge Gemini-2.5-Flash 95.4(± 1.6)95.4\,($\pm\,$1.6) 91.8(± 2.3)91.8\,($\pm\,$2.3) 98.3(± 1.2)98.3\,($\pm\,$1.2) 94.2(± 2.5)94.2\,($\pm\,$2.5) 95.1(± 2.4)95.1\,($\pm\,$2.4)
SurveyG GPT-4o-mini 96.2(± 1.4)96.2\,($\pm\,$1.4) 89.4(± 2.6)89.4\,($\pm\,$2.6) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{98.6}}\,(\pm\,1.0)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{95.6}}\,(\pm\,2.1)$ 96.2(± 2.3)96.2\,($\pm\,$2.3)
SurveyG Gemini-2.5-Flash $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}96.5}\,(\pm\,1.3)}$ $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}90.1}\,(\pm\,2.3)}$ 98.4(± 1.1)98.4\,($\pm\,$1.1) 95.2(± 2.3)95.2\,($\pm\,$2.3) $\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}96.7}\,(\pm\,2.1)}$
Table 7. Backbone comparison: SurveyForge and SurveyG with GPT-4o-mini vs. Gemini-2.5-Flash (mean (± std) over 10 trials per topic). Each LLM evaluates both models with different backbones across content quality dimensions.
Table 7 presents a systematic comparison of backbone models to isolate framework contributions from LLM capability effects. We evaluate both SurveyForge (current SOTA) and SurveyG with two different backbones: GPT-4o-mini and Gemini-2.5-Flash. The results demonstrate that backbone selection has minimal impact on relative performance rankings. For SurveyForge, switching backbones yields marginal differences (average $\pm$0.5-0.8 points across metrics), while SurveyG shows similarly stable performance (average variation <<0.6 points). Critically, SurveyG maintains consistent superiority over SurveyForge regardless of backbone choice: under Claude’s evaluation, SurveyG outperforms SurveyForge by +5.0 to +6.3 points in Coverage and +7.7 to +9.4 points in Structure across both backbones. This pattern holds across all four judges, with SurveyG achieving higher scores in 18 out of 20 model-metric combinations (90%). The consistency of performance gaps across different backbones confirms that SurveyG’s improvements stem from its hierarchical citation graph and multi-level summarization framework rather than simply leveraging more capable foundation models. These findings validate that our architectural innovations, not backbone selection, drive the observed quality improvements in automated survey generation.
Appendix C Experimental Setting
C.1. Baseline Systems
AutoSurvey (Wang et al., ) represents an early multi-stage framework for automated survey generation that combines literature retrieval with structured content synthesis. SurveyX (Liang et al., 2025) enhances AutoSurvey through structured knowledge extraction and outline optimization, introducing refinement mechanisms for improved coherence. SurveyForge (Yan et al., 2025) leverages human-written survey papers from related domains as prior knowledge for heuristic outline generation, guided by a memory-driven scholar navigation agent that retrieves high-quality references for composing new surveys.
C.2. Survey Topics and Ground Truth Construction
We compiled ten representative survey topics covering diverse research areas in machine learning and natural language processing, as summarized in Table 6. Each topic reflects an active line of inquiry, providing a strong foundation for evaluating automated survey generation.
C.2.1. Expert Recruitment and Qualifications
Our 20 domain experts comprised:
  • 12 CS Ph.D. students from QS 5-star universities with 3-5 years of research experience
  • 8 senior AI research engineers from leading tech companies with 2+ publications in top-tier venues (NeurIPS, ICML, ICLR, ACL, EMNLP, KDD, etc.)
Each expert was matched to topics within their domain expertise, ensuring a qualified assessment.
C.2.2. Three-Stage Ground Truth Selection Protocol
Stage 1: Topic Selection. We identified ten topics ensuring coverage across diverse AI/ML research areas: (a) foundational architectures (Visual Transformer, Graph Neural Networks, Generative Adversarial Networks), (b) learning paradigms (Deep Meta-Learning, Reinforcement Learning methods), (c) knowledge representation and retrieval (Knowledge Graph Embedding, Retrieval-Augmented Generation), and (d) reliability and robustness (Hallucination in LLMs, Generalized Out-of-Distribution Detection). This selection spans established methodologies, emerging challenges, and cross-domain techniques, ensuring comprehensive evaluation across theoretical foundations, architectural innovations, and practical applications. As shown in Table 6, selected ground-truth surveys are highly cited (mean: 484 citations), indicating strong community recognition.
Stage 2: Survey Selection Criteria. For each topic, 2-3 domain-matched experts independently evaluated candidate surveys from SurGE using four weighted criteria:
  1. (1)
    Comprehensive Coverage (40%): Survey thoroughly discusses field evolution, seminal works, major methodological branches, and current trends
  2. (2)
    Structural Clarity (25%): Well-organized sections with clear taxonomy, logical flow, and effective use of figures/tables
  3. (3)
    Recency (20%): Published within the last 5 years (2019-2024) to reflect the current research landscape
  4. (4)
    Citation Impact (15%): Highly-cited work (top 20% within topic area) indicating community recognition
Experts scored each criterion 0-10, with surveys scoring $\geq$ 7.0 average advancing to consensus discussion. Inter-annotator agreement (Fleiss’s $\kappa$) averaged 0.73 across topics, indicating substantial agreement.
Stage 3: Reference Curation. Selected experts curated 30-50 essential papers per topic, categorized as: foundational works (seminal papers establishing the field), methodological milestones (key technical innovations), and recent advances (papers from the last 2 years). This reference set serves as ground truth for coverage evaluation.
C.3. Evaluation Metrics
C.3.1. Outline Quality
Following (Yan et al., 2025), we evaluate outline structure, logical coherence, topic coverage, and hierarchical organization on a 0-100 scale. The evaluation prompt assesses whether the outline provides a clear roadmap for comprehensive survey development.
C.3.2. Content Quality Metrics
A well-structured outline is essential for maintaining clarity and organization. Full paper evaluation serves as a comprehensive benchmark to assess academic rigor and practical utility. Following (Wang et al., ; Liang et al., 2025; Yan et al., 2025; Su et al., 2025), we assess five metrics on a 0-100 scale:
Coverage measures how thoroughly the survey captures major concepts, foundational works, and emerging trends within the research area.
Structure examines logical organization, section coherence, and taxonomy quality, ensuring readers can navigate the field systematically.
Relevance assesses alignment of content with the target research topic, penalizing off-topic discussions or tangential references.
Synthesis evaluates integration of information from multiple sources into a cohesive, non-redundant narrative that reveals cross-paper insights.
Critical Analysis reflects the survey’s ability to identify methodological gaps, highlight comparative trends, and articulate open research challenges.
Each metric is scored independently by both LLM judges and human experts, with final scores averaged across all evaluators.
C.3.3. Citation Quality Metrics
Following (Wang et al., ; Gao et al., 2023), we measure citation accuracy and contextual relevance. We extract factual claims from generated surveys and verify support using a Natural Language Inference (NLI) model (DeBERTa-v3-large fine-tuned on ANLI).
Citation Recall: Proportion of claims correctly supported by valid references, measuring whether surveys provide adequate citation support.
Citation Precision: Proportion of cited references that truly substantiate their associated claims, measuring citation accuracy.
Citation F1: Harmonic mean of Recall and Precision, providing a balanced measure of citation quality.
For each survey, we sample 100 claim-citation pairs and compute metrics across all samples.
Appendix D Prompt Templates
@@
D.1. Prompt to generate structured outline
We provide a short version of the prompt template (Figure 8) that instructs the model to construct a coherent, hierarchical outline that captures the logical flow of a literature review topic before detailed writing begins.
@@ D.1. Prompt to generate structured outline
D.2. Prompt to evaluate structured outline
@@
D.3. Prompt to generate subsections
@@
Figure 10. Generate Subsection Prompt
D.4. Improve Section Quality
As shown in Figure 11, this prompt systematically assesses literature review sections across multiple dimensions, such as content coverage, synthesis, and critical analysis, while offering actionable feedback and retrieval suggestions for refinement.
@@ D.4. Improve Section Quality
Appendix E Computational Efficiency
We analyze computational overhead for the hierarchical citation graph operations on a standard workstation (1 NVIDIA RTX 4090, 24GB VRAM):
Paper Retrieval and Graph Construction: For a typical query retrieving 1,500 candidate papers, the process includes: (1) keyword expansion via LLM (2-3 seconds, \$0.02), (2) Semantic Scholar API calls (15-20 seconds, free), (3) citation network construction (8-12 seconds), and (4) embedding generation using Deberta (45-60 seconds on CPU, negligible GPU cost if available). Total: $\sim$70-95 seconds, \$0.02.
Hierarchical Layer Assignment: Organizing 300 papers into Foundation/Development/Frontier layers using citation patterns and temporal analysis requires 5-8 seconds (pure algorithmic processing, no API calls).
Multi-Aspect Summarization: Generating layer-wise and cross-layer summaries involves: (1) horizontal clustering within layers (10-15 seconds for Leiden algorithm), (2) vertical path traversal (3-5 seconds), and (3) LLM-based summarization of $\sim$15-20 paper groups (\$0.80-\$1.20 depending on summary length). Total: $\sim$15-25 seconds computation, \$0.80-\$1.20 LLM cost.
End-to-End Latency: The complete pipeline for generating one survey consists of: (1) Crawled data: 20-25 minutes, (2) Graph construction and traversal: 90-128 seconds, (2) Outline generation: 120-180 seconds, (3) Full survey writing: 5-10 minutes (parallel generation of 8-12 sections with evaluation loops). Total end-to-end latency: 28-38 minutes per survey, with the majority of time spent on crawled data rather than graph operations or content generation.
Appendix F Case studies
@@
Figure 12. Case studies about the result of the generated subsection.